OpenStudy (anonymous):

A circle that has a radius of 6 units has a circumference of 37.68 units. What is the circumference of a circle that has a radius of 24 units? 602.88 units 150.72 units 904.32 units 678.24 units

OpenStudy (anonymous):

unit analysis radius = 6 units, circumference = 37.68 units circumference/radius is unitless; unitless values are constant circumference/radius = 37.68units/6units = 37.68/6 = 6.28 radius = 24 units, circumference = ? units circumference/radius = ?/24 = 6.28 find ?

OpenStudy (anonymous):

602.88 @eashy

OpenStudy (anonymous):

? = 6.28*24 = 150.72

OpenStudy (anonymous):

oh I get it